0.98 lbm/s of Nitrogen initially at 390 psia, 9 ft3/min, 299°F are cooled isometricall to 118°F in an internally reversible manner. Compute P2, ⌠pdv, -⌠vdp, ▲U, ▲H, Q and ▲S for a nonflow process. (This is the exact question in the book Sir/Ma'am)


1
Expert's answer
2021-02-25T04:09:51-0500

⌠pdv = nR(T2 - T1)/V = 0.98 × 8.314 ×{ 299-118)/390 =

1474.74/390

= 3.78 N/m